An object moves uniformly around a circular path of radius 23.0 cm, making one complete revolution every 1.85 s.

(a) What is the translational speed of the object?
____m/s?

(b) What is the frequency of motion in hertz?
____Hz?

(c) What is the angular speed of the object?
___rad/s ?
Answer #1

a) v = 2*pi*0.23/1.85 = 0.78075 m/s
b) f = 1/T = 0.54054 Hz
c) w = v/r = 3.3945 rad/s